A solution is prepared by dissolving 0.0034 g Ca(OH)2 in enough water to prepare a 250 mL solution. what is the pH of the solution if Ca(OH)2 completely dissociates producing 2 OH for every 1 Ca
Answer #1

Molaro mass of Ca(OH) = 74.092 g/mol Numbers of moles of CaCOH & Weight . molars mass 0.0034 g 74.092 g farol - 4.58× 10-5 mo.: pH = 14-POH = 14 -3.43 = 10.57